# On the complementation of spaces of $\mathcal{I}$-null sequences

### v1

@@ 5. When is $c_{0,\mathcal{I}}$ complemented in $\ell_{\infty}$?
Let $\{K_{n}\,\colon\,n\in\mathbb{N}\}$ be a partition of $\mathbb{N}$ and let $\mathcal{I}_{n}$ be an ideal on $K_{n}$ for each $n\in\mathbb{N}$. Suppose that $c_{0,\mathcal{I}_{n}}$ is complemented in $\ell_{\infty}(K_{n})$ for each $n\in\mathbb{N}$. Then, the ideal $\mathcal{J}=\bigoplus_{n\in\mathbb{N}}\mathcal{I}_{n}$ is complemented. In particular, if $\mathcal{I}$ is a strongly $\omega$-maximal ideal, then $\mathcal{I}$ is complemented.
@@
By the aforementioned Lindenstrauss’s theorem, $c_{0,\mathcal{I}_{m}}$ is isomorphic to $\ell_{\infty}(K_{m})$ for each $m\in\mathbb{N}$. From [23, Theorem 5.2] it follows that
@@
Again, by Lindenstrauss’s theorem, we conclude that $c_{0,\mathcal{J}}$ is complemented in $\ell_{\infty}$.
Suppose $\mathcal{I}$ is strongly $\omega$-maximal. By Proposition 3.1, there is a partition $\{K_{n}\}$ of $\mathbb{N}$ such that $\mathcal{I}=\bigoplus_{n\in\mathbb{N}}(\mathcal{I}_{n}\restriction K_{n})$ and $\mathcal{I}_{n}\restriction K_{n}$ is maximal on $K_{n}$ for each $n$. Then, the previous argument shows that $c_{0,\mathcal{I}}$ is isomorphic to $\ell_{\infty}$. ∎
@@
It follows from Theorem 5.2 that if $\mathcal{I}$ is complemented, then $\mathcal{I}^{\omega}$ is complemented. Conversely, suppose that $\mathcal{I}^{\omega}$ is complemented. By [23, Theorem 5.2] we have $c_{0,\mathcal{I}^{\omega}}\cong\ell_{\infty}(c_{0,\mathcal{I}})$. Consequently, $c_{0,\mathcal{I}}$ is isomorphic to a complemented subspace of $c_{0,\mathcal{I}^{\omega}}$. On the other hand, notice that $c_{0,\mathcal{I}^{\omega}}\sim\ell_{\infty}$ by Lindenstrauss’s theorem. Thus, $c_{0,\mathcal{I}}$ is isomorphic to a complemented subspace of $\ell_{\infty}$. Therefore, $\mathcal{I}$ is complemented. ∎
Proposition 5.4.
@@
Despite these results, it remains unknown whether the intersection of a countable family of complemented ideals is itself complemented. Even in the case of two complemented ideals, $\mathcal{I}$ and $\mathcal{J}$, we do not know whether $\mathcal{I}\cap\mathcal{J}$ is complemented.
@@
Let $\mathcal{I}$ be an ideal on $\mathbb{N}$ and let $A$ and $B$ be infinite subsets of $\mathbb{N}$ with $A\subseteq B$ and $A\in\mathcal{I}^{+}$. If $\mathcal{I}\restriction B$ is complemented in $B$, then $\mathcal{I}\restriction A$ is complemented in $A$. In particular, if $\mathcal{I}$ is complemented, then $\mathcal{I}\restriction A$ is complemented in $A$ for every $A\in\mathcal{I}^{+}$.
@@
Now we present a more informative proof of Theorem 5.2. But first we need an auxiliary result. Recall that if ${\bf x}=(x_{n})\in\ell_{\infty}$, its support is defined as $\mathrm{supp}({\bf x})\coloneqq\{n\in\mathbb{N}\,\colon\,x_{n}\neq 0\}$. Let $F\subseteq\mathbb{N}$, and suppose that $({\bf x}_{n})_{n\in F}$ is a sequence in $\ell_{\infty}$ such that $\mathrm{supp}({\bf x}_{n})\cap\mathrm{supp}({\bf x}_{n})=\emptyset$ for all $m\neq n$. Then, the sum $\sum_{n\in F}{\bf x}_{n}$ is the sequence in $\ell_{\infty}$ whose $m$-th coordinate is ${\bf x}_{n}(m)$ whenever $m\in\mathrm{supp}({\bf x}_{n})$ for some $n\in F$, and 0 otherwise.
@@
    The family =$\mathcal{B}=\{B\subseteq\mathbb{N}\,\colon\,T(B)=B\}$ is closed under arbitrary intersections.
@@
Now define $Q\colon\ell_{\infty}\to\ell_{\infty}$ by $Q=Id-P$, that is,
@@
Remark 5.10.
Property (4) of Theorem 5.8 can be restated as follows: For any pair $A,B$ of subsets of $\mathbb{N}$ and $n\in\mathbb{N}$,
We will show that, under some extra condition, the complementation of $\mathcal{I}$ implies that $\mathcal{I}$ is $\omega$-maximal. For that end, we need a key lemma whose argument is motivated by the proof of the non-complementation of $c_{0}$ in $\ell_{\infty}$ due to Whitley (see [7, Theorem 5.6] or [1, Theorem 2.5.4]). Indeed, we will show a more general fact: $c_{0}$ is not complemented in $c_{0,\mathcal{J}}$ for any ideal $\mathcal{J}\supsetneq{\sf Fin}$. A similar result appears in [18, Corollary 1.5] about the space $c$ of convergent sequences.
Lemma 5.11.
Let $\mathcal{I}\subsetneq\mathcal{J}$ be ideals on $\mathbb{N}$ such that $c_{0,\mathcal{I}}$ is complemented in $c_{0,\mathcal{J}}$. Then, there exists a collection $\{\mathcal{A}_{n,k}\colon\,n,k\in\mathbb{N}\}$ of subsets of $2^{\mathbb{N}}$ satisfying:
  1. (1)
  2. (2)
    $\mathcal{A}_{n,k}\subseteq\mathcal{A}_{n,k+1}$ for all $n,k\in\mathbb{N}$.
  3. (3)
  4. (4)
  5. (5)
Proof.
Let $P\colon c_{0,\mathcal{J}}\to c_{0,\mathcal{I}}$ be a projection onto $c_{0,\mathcal{I}}$ and $Q=Id-P$. For each $n,k\in\mathbb{N}$, define
$\mathcal{A}_{n,k}=\{A\in\mathcal{J}\,\colon\,|Q(\chi_{A})_{n}|\geq 1/k\}.$
Since =$\mathcal{I}=\{A\in\mathcal{J}\,\colon\,Q(\chi_{A})={\bf 0}\}$, we have
$\mathcal{J}\cap\mathcal{I}^{+}$ =$=\{A\in\mathcal{J}\,\colon\,Q(\chi_{A})\neq{\bf 0}\}$
$=\bigcup_{n\in\mathbb{N}}\mathcal{A}_{n}.$
$=\sum_{j=1}^{m}a_{j}Q(\chi_{F_{j}})_{n}$
$\leq\left\|\sum_{j=1}^{m}a_{j}Q(\chi_{F_{j}})\right\|$
$\leq\|Q\|\left\|\sum_{j=1}^{m}a_{j}\chi_{F_{j}}\right\|=\|Q\|.$
Hence, $m\leq k\|Q\|$.
$|Q(\chi_{A})_{n}|=|Q(\chi_{A\setminus B\cap C})_{n}|=|Q(\chi_{B})_{n}+Q(\chi_{C})_{n}|\leq 1/k,$
which is absurd. ∎
Theorem 5.12.
Let $\mathcal{I}\subsetneq\mathcal{J}$ be proper ideals on $\mathbb{N}$ such that $c_{0,\mathcal{I}}$ is complemented in $c_{0,\mathcal{J}}$. Then $\mathfrak{ad}_{\mathcal{I}}(\mathcal{J}\cap\mathcal{I}^{+})\leq\aleph_{0}$.
Proof.
For $k,n\in\mathbb{N}$, let $\mathcal{A}_{n,k}$ be as in Lemma 5.11. Then, we have $\mathcal{J}\cap\mathcal{I}^{+}=\bigcup_{n,k}\mathcal{A}_{n,k}$ and $\mathfrak{ad}_{\mathcal{I}}(\mathcal{A}_{n,k})<\aleph_{0}$ for each $n,k$. Suppose $\mathcal{B}\subseteq\mathcal{J}\cap\mathcal{I}^{+}$ is an $\mathcal{I}$-AD family. Then, for every $n,k\in\mathbb{N}$, $\mathcal{B}\cap\mathcal{A}_{n,k}$ is finite. Thus, $\mathcal{B}$ is countable, and consequently $\mathfrak{ad}_{\mathcal{I}}(\mathcal{J}\cap\mathcal{I}^{+})\leq\aleph_{0}$. ∎
The next observation follows immediately from Lemma 5.1.
Corollary 5.13.
Let $\mathcal{I}\subsetneq\mathcal{J}$ be proper ideals on $\mathbb{N}$ such that $\mathcal{I}\restriction A$ is meager as a subset of $2^{A}$ for some infinite set $A\in\mathcal{J}$. Then $\mathfrak{ad}_{\mathcal{I}}(\mathcal{J}\cap\mathcal{I}^{+})>\aleph_{0}$.
From Theorem 5.12 we get the following.
Corollary 5.14.
Remark 5.15.
Proposition 5.16.
Proof.
However, it is unclear whether $c_{0,\mathcal{I}}$ contains a complemented copy of $c_{0}$ for any meager ideal $\mathcal{I}$.
The next corollary is due to Kania [14, Theorem A].
Corollary 5.17.
Let $\mathcal{I}$ be an ideal on $\mathbb{N}$. If $\mathcal{I}$ is complemented, then $\mathfrak{ad}(\mathcal{I}^{+})\leq\aleph_{0}$.
Now we prove, with an extra assumption, that the complementation of $c_{0,\mathcal{I}}$ implies that $\mathcal{I}$ is $\omega$-maximal.
Theorem 5.18.
Let $\mathcal{I}$ be a proper ideal on $\mathbb{N}$. Suppose $c_{0,\mathcal{I}}$ is complemented in $\ell_{\infty}$ by a projection map $P\colon\ell_{\infty}\to c_{0,\mathcal{I}}$ such that $Q:=Id-P$ satisfies (5.2). Then $\mathcal{I}$ is $\omega$-maximal.
Proof.
Suppose $P\colon\ell_{\infty}\to c_{0,\mathcal{I}}$ is a projection map as in the hypothesis. For each $n\in\mathbb{N}$ consider the family
$\mathcal{A}_{n}=\{A\subseteq\mathbb{N}:\;Q(\chi_{A})_{n}\neq 0\}$
By (4) of Lemma 5.11, $\mathcal{I}=\bigcap_{n}\mathcal{I}_{n}=\bigcap_{n\in S}\mathcal{I}_{n}$. It remains to show that each $\mathcal{I}_{n}$ is maximal for each $n\in S$. If not, let $B\subseteq\mathbb{N}$ be such that $B\not\in\mathcal{I}_{n}$ and $\mathbb{N}\setminus B\not\in\mathcal{I}_{n}$. Then there are $C_{0},C_{1}\in\mathcal{A}_{n}$ such that $C_{0}\subseteq B$ and $C_{1}\subseteq\mathbb{N}\setminus B$. So, $Q(\chi_{C_{0}\cap C_{1}})={\bf 0}$. On the other hand, by (5.2) we have $Q(\chi_{C_{0}\cap C_{1}})_{n}\neq 0$, an absurd. ∎
Remark 5.19.
It is not difficult to verify that $\mathcal{I}_{n}$ is maximal if and only if Property (5.2) holds for that $n$. We do not know whether the assumption of Property (5.2) can be removed.
Theorem 5.18 raises the following question: Is there an $\omega$-maximal ideal that is not complemented? In other words, is being complemented equivalent to being $\omega$-maximal? A natural candidate for a negative answer to this question is the $\omega$-maximal ideal constructed in Example 3.8.
@@
  • The family =$\mathcal{B}=\{B\subseteq\mathbb{N}\,\colon\,T(B)=B\}$ is closed under arbitrary intersections.

### v2

@@ 5. Structural properties of complemented ideals
Let $\{K_{n}\,\colon\,n\in\mathbb{N}\}$ be a partition of $\mathbb{N}$ and let $\mathcal{I}_{n}$ be an ideal on $K_{n}$ for each $n\in\mathbb{N}$. Suppose that $c_{0,\mathcal{I}_{n}}$ is complemented in $\ell_{\infty}(K_{n})$ for each $n\in\mathbb{N}$. Then, the ideal $\mathcal{J}=\bigoplus_{n\in\mathbb{N}}\mathcal{I}_{n}$ is complemented.
@@
By the aforementioned Lindenstrauss’s theorem, $c_{0,\mathcal{I}_{m}}$ is isomorphic to $\ell_{\infty}(K_{m})$ for each $m\in\mathbb{N}$. From [25, Theorem 5.2] and Remark 5.4 it follows that
@@
Again, by Lindenstrauss’s theorem, we conclude that $c_{0,\mathcal{J}}$ is complemented in $\ell_{\infty}$. ∎
@@
It follows from Theorem 5.5 that if $\mathcal{I}$ is complemented, then $\mathcal{I}^{\omega}$ is complemented. Conversely, suppose that $\mathcal{I}^{\omega}$ is complemented. By [25, Theorem 5.2] we have $c_{0,\mathcal{I}^{\omega}}\cong\ell_{\infty}(c_{0,\mathcal{I}})$. Consequently, $c_{0,\mathcal{I}}$ is isomorphic to a complemented subspace of $c_{0,\mathcal{I}^{\omega}}$. On the other hand, by Lindenstrauss’s theorem, $c_{0,\mathcal{I}^{\omega}}\sim\ell_{\infty}$. Hence, $c_{0,\mathcal{I}}$ is isomorphic to a complemented subspace of $\ell_{\infty}$. Therefore, by Lindenstrauss’s theorem, $c_{0,\mathcal{I}}$ is isomorphic to $\ell_{\infty}$, and thus $\mathcal{I}$ is complemented. ∎
Theorem 5.7.
Let $\{\mathcal{I}_{n}\,\colon\,n\in\mathbb{N}\}$ be a collection of complemented ideals. Then, $\mathcal{I}=\bigcap_{n\in\mathbb{N}}\mathcal{I}_{n}$ is complemented. In particular, every $\omega$-maximal ideal is complemented.
@@
Let $A\colon\ell_{\infty}\to\ell_{\infty}((\ell_{\infty}/c_{0,\mathcal{I}_{n}})_{n\in\mathbb{N}})$ be defined by $A({\bf x})=({\bf x}+c_{0,\mathcal{I}_{n}})$, $\bf x$$\in\ell_{\infty}$. Then $A$ is linear and continuous, and $\ker A=c_{0,\mathcal{I}}$. Hence, $\ell_{\infty}/c_{0,\mathcal{I}}$ is isomorphic to a subspace of $\ell_{\infty}((\ell_{\infty}/c_{0,\mathcal{I}_{n}})_{n\in\mathbb{N}})$.
On the other hand, since $\mathcal{I}_{n}$ is complemented for each $n\in\mathbb{N}$, it follows from Theorem 5.2, that $\ell_{\infty}/c_{0,\mathcal{I}_{n}}$ is isomorphic to $\ell_{\infty}$ for every $n\in\mathbb{N}$. Therefore, $\ell_{\infty}((\ell_{\infty}/c_{0,\mathcal{I}_{n}})_{n\in\mathbb{N}})\sim\ell_{\infty}(\ell_{\infty})\sim\ell_{\infty}$. Consequently, $\ell_{\infty}/c_{0,\mathcal{I}}$ is isomorphic to a subspace of $\ell_{\infty}$ and $\mathcal{I}$ is complemented. ∎
@@
Example 2 in [13] shows that the second part of Theorem 5.7 is not an equivalence.
@@
Let $\mathcal{I}$ be an ideal on $\mathbb{N}$, and let $A$ and $B$ be infinite subsets of $\mathbb{N}$ with $A\subseteq B$ and $A\in\mathcal{I}^{+}$. If $\mathcal{I}\restriction B$ is complemented in $B$, then $\mathcal{I}\restriction A$ is complemented in $A$. In particular, if $\mathcal{I}$ is complemented, then $\mathcal{I}\restriction A$ is complemented in $A$ for every $A\in\mathcal{I}^{+}$.
@@
We will show that, under an additional condition, the complementation of $\mathcal{I}$ is equivalent to $\mathcal{I}$ being $\omega$-maximal. For that end, we need a key lemma whose argument is motivated by the proof of the non-complementation of $c_{0}$ in $\ell_{\infty}$ due to Whitley (see [7, Theorem 5.6] or [1, Theorem 2.5.4]).
Lemma 5.10.
Let $\mathcal{I}\subsetneq\mathcal{J}$ be ideals on $\mathbb{N}$. Suppose that there is a linear bounded operator $Q\colon c_{0,\mathcal{J}}\to c_{0,\mathcal{J}}$ such that $\ker Q=c_{0,\mathcal{I}}$. Then, there exists a collection $\{\mathcal{A}_{n,k}\colon\,n,k\in\mathbb{N}\}$ of subsets of $2^{\mathbb{N}}$ satisfying:
  1. (1)
  2. (2)
    $\mathcal{A}_{n,k}\subseteq\mathcal{A}_{n,k+1}$ for all $n,k\in\mathbb{N}$.
  3. (3)
  4. (4)
  5. (5)
Proof.
For each $n,k\in\mathbb{N}$, define
$\mathcal{A}_{n,k}=\{A\in\mathcal{J}\,\colon\,|Q(\chi_{A})_{n}|\geq 1/k\}.$
Since $\mathcal{I}=\{A\in\mathcal{J}\,\colon\,Q(\chi_{A})={\bf 0}\}$, we have
$\mathcal{J}\cap\mathcal{I}^{+}$ $=\{A\in\mathcal{J}\,\colon\,Q(\chi_{A})\neq{\bf 0}\}$
$=\bigcup_{n\in\mathbb{N}}\mathcal{A}_{n}.$
$=\sum_{j=1}^{m}a_{j}Q(\chi_{F_{j}})_{n}$
$\leq\left\|\sum_{j=1}^{m}a_{j}Q(\chi_{F_{j}})\right\|$
$\leq\|Q\|\left\|\sum_{j=1}^{m}a_{j}\chi_{F_{j}}\right\|=\|Q\|.$
Hence, $m\leq k\|Q\|$.
$|Q(\chi_{A})_{n}|=|Q(\chi_{A\setminus B\cap C})_{n}|=|Q(\chi_{B})_{n}+Q(\chi_{C})_{n}|<1/k,$
which is absurd. ∎
Now we prove, with an extra assumption, that the complementation of $c_{0,\mathcal{I}}$ is equivalent to $\mathcal{I}$ being $\omega$-maximal.
Theorem 5.11.
Let $\mathcal{I}$ be a proper ideal on $\mathbb{N}$. Then, the following statements are equivalent:
  1. (1)
    $\mathcal{I}$ is $\omega$-maximal.
  2. (2)
    There is a Banach lattice isomorphism from $\ell_{\infty}/c_{0,\mathcal{I}}$ into $\ell_{\infty}$.
  3. (3)
    There exists a bounded linear operator $Q\colon\ell_{\infty}\to\ell_{\infty}$ with $\ker Q=c_{0,\mathcal{I}}$ that verifies the next property: for every pair of subsets $A,B\subseteq\mathbb{N}$ and every $n\in\mathbb{N}$,
Proof.
$(1)\Rightarrow(2)$ Notice that $\ell_{\infty}/c_{0,\mathcal{I}_{n}}$ is Banach lattice isomorphic to $\mathbb{R}$ for each $n\in\mathbb{N}$ via the map ${\bf x}+c_{0,\mathcal{I}_{n}}\mapsto\mathcal{I}_{n}-\lim{\bf x}$. Hence, $\ell_{\infty}((\ell_{\infty}/c_{0,\mathcal{I}_{n}})_{n\in\mathbb{N}})$ is Banach lattice isomorphic to $\ell_{\infty}$. Now the map $L\colon\ell_{\infty}/c_{0,\mathcal{I}}\to\ell_{\infty}((\ell_{\infty}/c_{0,\mathcal{I}_{n}})_{n\in\mathbb{N}})$, ${\bf x}+c_{0,\mathcal{I}}\mapsto({\bf x}+c_{0,\mathcal{I}_{n}})_{n\in\mathbb{N}}$ is a Banach lattice embedding. Thus, there is a Banach lattice isomorphism from $\ell_{\infty}/c_{0,\mathcal{I}}$ into $\ell_{\infty}$.
$(2)\Rightarrow(3)$ Let $T\colon\ell_{\infty}/c_{0,\mathcal{I}}\to\ell_{\infty}$ be a Banach lattice embedding. Define $Q\colon\ell_{\infty}\to\ell_{\infty}$ by $Q({\bf x})=T({\bf x}+c_{0,\mathcal{I}})$ for each $\bf x$$\in\ell_{\infty}$. It is clear that $Q$ is linear and bounded. Since $T$ is injective, we have:
$Q({\bf x})={\bf 0}\Longleftrightarrow T({\bf x}+c_{0,\mathcal{I}})={\bf 0}\Longleftrightarrow{\bf x}\in c_{0,\mathcal{I}}.$
Thus, $\ker Q=c_{0,\mathcal{I}}$.
We will verify (5.1). Let $n\in\mathbb{N}$ and $A,B\subseteq\mathbb{N}$ be such that $Q(\chi_{A})_{n}\neq 0$ and $Q(\chi_{B})_{n}\neq 0$. Observe that $Q(\chi_{A})=T(\chi_{A}+c_{0,\mathcal{I}})\geq{\bf 0}$ and $Q(\chi_{B})=T(\chi_{B}+c_{0,\mathcal{I}})\geq{\bf 0}$. Hence, $Q(\chi_{A})_{n}=T(\chi_{A}+c_{0,\mathcal{I}})_{n}>0$ and $Q(\chi_{B})_{n}=T(\chi_{B}+c_{0,\mathcal{I}})_{n}>0$. From the fact $(\chi_{A}+c_{0,\mathcal{I}})\wedge(\chi_{B}+c_{0,\mathcal{I}})=\chi_{A\cap B}+c_{0,\mathcal{I}}$, it follows that $T(\chi_{A\cap B}+c_{0,\mathcal{I}})=T(\chi_{A}+c_{0,\mathcal{I}})\wedge T(\chi_{A}+c_{0,\mathcal{I}})$ and
$Q(\chi_{A\cap B})_{n}=T(\chi_{A\cap B}+c_{0,\mathcal{I}})_{n}=T(\chi_{A}+c_{0,\mathcal{I}})_{n}\wedge T(\chi_{A}+c_{0,\mathcal{I}})_{n}>0.$
Thus, $Q(\chi_{A\cap B})_{n}\neq 0$.
$(3)\Rightarrow(1)$ Let $Q\colon\ell_{\infty}\to c_{0,\mathcal{I}}$ be a bounded linear operator satisfying (5.1). For each $n\in\mathbb{N}$, consider the family
$\mathcal{A}_{n}=\{A\subseteq\mathbb{N}:\;Q(\chi_{A})_{n}\neq 0\}$
By (4) of Lemma 5.10, $\mathcal{I}=\bigcap_{n}\mathcal{I}_{n}=\bigcap_{n\in S}\mathcal{I}_{n}$. It remains to show that each $\mathcal{I}_{n}$ is maximal for each $n\in S$. If not, let $B\subseteq\mathbb{N}$ be such that $B\not\in\mathcal{I}_{n}$ and $\mathbb{N}\setminus B\not\in\mathcal{I}_{n}$. Then there are $C_{0},C_{1}\in\mathcal{A}_{n}$ such that $C_{0}\subseteq B$ and $C_{1}\subseteq\mathbb{N}\setminus B$. So, $Q(\chi_{C_{0}\cap C_{1}})={\bf 0}$. On the other hand, by (5.1) we have $Q(\chi_{C_{0}\cap C_{1}})_{n}\neq 0$, an absurd. This completes the proof. ∎
Remark 5.12.
The Property (5.1) cannot be omitted. Indeed, by Corollary 5.3, the existence of a bounded linear operator $Q$ from $\ell_{\infty}$ to $\ell_{\infty}$ with $\ker Q=c_{0,\mathcal{I}}$ implies that $\mathcal{I}$ is complemented. However, [13, Example 2] shows a complemented ideal which is not $\omega$-maximal.
@@
Observe that if $\mathcal{I}$ is a strongly $\omega$-maximal ideal, then $\mathcal{I}$ is complemented by Theorem 5.7. Our next result provides an alternative proof by constructing an explicit projection. To this end, we first establish an auxiliary result. Recall that if ${\bf x}=(x_{n})\in\ell_{\infty}$, its support is defined as $\mathrm{supp}({\bf x})\coloneqq\{n\in\mathbb{N}\,\colon\,x_{n}\neq 0\}$. Let $F\subseteq\mathbb{N}$, and suppose that $({\bf x}_{n})_{n\in F}$ is a sequence in $\ell_{\infty}$ such that $\mathrm{supp}({\bf x}_{n})\cap\mathrm{supp}({\bf x}_{n})=\emptyset$ for all $m\neq n$. Then, the sum $\sum_{n\in F}{\bf x}_{n}$ is the sequence in $\ell_{\infty}$ whose $m$-th coordinate is ${\bf x}_{n}(m)$ whenever $m\in\mathrm{supp}({\bf x}_{n})$ for some $n\in F$, and 0 otherwise. Observe that the previous sum is well defined since the supports are pairwise disjoint.
@@
Next result presents an explicit projection in the case of strongly $\omega$-maximal ideals.
@@
    The family $\mathcal{B}=\{B\subseteq\mathbb{N}\,\colon\,T(B)=B\}$ is closed under arbitrary intersections.
@@
Now define $Q\colon\ell_{\infty}\to\ell_{\infty}$ by Q=IdPQ=Id-P, that is,
@@
  • The family $\mathcal{B}=\{B\subseteq\mathbb{N}\,\colon\,T(B)=B\}$ is closed under arbitrary intersections.